# A Correspondence-Driven Approach for Bilevel Decision-making with Nonconvex Lower-Level Problems

### v1

@@ 6 Conclusion
Building on these results, we designed a biased projected SGD-based algorithm SCiNBiO with a cubic-regularized Newton lower-level solver, and provided convergence guarantees together with oracle complexity bounds for the upper level. Under the additional assumption that all degenerate stationary points are fold bifurcation points, we further obtained the lower-level oracle complexity of SCiNBiO.
Our work develops new modeling and algorithmic approaches for bilevel optimization with nonconvex lower levels under minimal assumptions, offering practical solution methods together with theoretical insights and provable performance guarantees. Future work will aim to extend these results to settings with stochastic objectives or different lower-level solvers.
@@
Appendix A Experimental Setup and Full Results
A.1 Numerical Example
@@ Appendix B Measure-Zero Discontinuities of Correspondence-Driven Solutions under Gradient Descent
Suppose $g(x,y)$ is a $C^{2}$ function such that it is coercive and semi-algebraic in $y$ for any $x$, $\mathcal{M}$ is gradient descent method, and $0<\eta<1/\overline{L}_{g}$. We uniformly choose $b$ from $\nu,\nu]^{m}$, where $\nu>0$ can be any constant, and perturb the initial point $y_{0}$ by setting it to $y_{0}+b$. Then, with probability one, the set of points in $\mathcal{X}\setminus\widetilde{\mathcal{X}}$ at which $y^{\text{cd}}(x)$ is discontinuous has measure zero, where $y^{\mathrm{cd}}(x)$ is defined under the perturbed function $\tilde{g}(x,y)$.
@@
    Consider the map $F_{x}:y\mapsto y-\eta\nabla_{y}g(x,y)$, then $F_{x}(y)$ is a diffeomorphism from $\mathbb{R}^{m}\to\mathbb{R}^{m}$;
@@
Next, we prove that $F_{x}$ is subjective. For any $y_{1}$, consider the following problem
@@
Detailed proof of Theorem B.1: Since $g(x,y)$ is a Morse function in $y$ for any $x\in\mathcal{X}\setminus\widetilde{\mathcal{X}}$, all stationary points are non-degenerate and isolated. As a result, the set of stationary points is countable and denoted by $\{p_{j}\}_{j\in J}$, where $J$ is the index set. Since $g(x,y)$ is coercive and semi-algebraic in $y$, for any initial point, the gradient descent method can generate a sequence converges to a stationary (Attouch et al., 2013), so $\mathbb{R}^{m}=\cup_{j\in J}\mathcal{W}^{s}(p_{j})$, where $\mathcal{W}^{s}(p_{j})$ is the stable manifold with respect to $p_{j}$. By the global version of the stable manifold theorem, $\mathcal{W}^{s}(p_{j})$ is homeomorphic to the open unit ball of dimension $n_{j}$, where $n_{j}$ is the number of positive eigenvalues of $\nabla^{2}_{yy}g(x,y)$ at $p_{j}$. Therefore, the stable manifold of a local minimum is an open $m$-dimensional submanifold of $\mathbb{R}^{m}$, whereas the stable manifolds of the other stationary points, denoted as $\mathcal{A}$, have lower dimension and thus measure zero. Note that we uniformly choose $b$ from $\nu,\nu]^{m}$ and perturb the initial point $y_{0}$ as $y_{0}+b$, it follows that the probability that $y_{0}+b$ lies in $\mathcal{A}$ is zero.
In the following discussion, we may assume without loss of generality that $\mathcal{X}/\widetilde{\mathcal{X}}$ is simply connected, i.e., it has only one connected component, as the following analysis can be repeated on each connected component.
By the implicit function theorem and Morse property of $g(x,y)$ in $y$, each critical point $p_{j}(x)$ corresponding to a solution of $\nabla_{y}g(x,y)=0$ varies continuously with $x$. From the point of view of Jacob Palis (1993, Page 165), we know that any stable manifold $\mathcal{W}^{s}(p_{j}(x)))$ with respect to the stationary point $p_{j}(x)$ continuously deforms in $x$ (in the sense that $\cup_{x\in\mathcal{X}\setminus{\widetilde{\mathcal{X}}}}(\{x\}\times W^{s}(p_{j}(x)))$ is a injectively immersed submanifold of $\mathcal{X}\setminus{\widetilde{\mathcal{X}}}\times\mathbb{R}^{m}$). Note that the stable manifold of a local minimizer $p_{j}(x)$ is an open $m$-dimensional submanifold of $\mathbb{R}^{m}$. It is not hard to say that if the initialization $y_{0}$ lies in the stable manifold of some local minimizer $p_{j}(x)$, i.e., gradient descent with step size $\eta$ will converge to $y^{\text{cd}}(x)=p_{j}(x)$, then under a small perturbation of $x$ to $x+\Delta x$, the initialization $y_{0}$ will still lie in the stable manifold of the corresponding minimizer $p_{j}(x+\Delta x)$. This implies that $y^{\text{cd}}(x)$ is continuous at this $x$. In summary, if $y_{0}$ lies in the stable manifold of a local minimizer, then $y^{\text{cd}}(x)$ is continuous at $x$. Discontinuity can therefore only occur when $y_{0}$ lies in the stable manifold of a non-minimizing stationary point, such as a saddle point or a local maximizer. Letting $\{p_{j}(x)\}_{j\in J^{\prime}}$ denote the collection of such saddle points and local maximizers where $J^{\prime}$ is a subset of $J$, we define the corresponding union of stable manifolds and the indicator function as follows:
$\mathcal{A}(x):=\bigcup_{j\in J^{\prime}}\mathcal{W}^{s}(p_{j}(x)),$
$I_{\mathcal{A}(x)}(y_{0},b)=\begin{cases}1&\text{ if }y_{0}+b\in\mathcal{A}(x)\\ 0&\text{ if }y_{0}+b\notin\mathcal{A}(x).\end{cases}$
Since $\mathbb{R}^{m}\setminus\mathcal{A}(x)$ is the union of stable manifolds of strict local minima (each with a positive definite Hessian), which are $m$-dimensional open submanifolds, it follows that $\mathcal{A}(x)$ is closed. Therefore, $I_{\mathcal{A}(x)}(y_{0},b)=1$ if and only if $D(x,b)=0$, where $D(x,b):=\mathrm{dist}(y_{0}+b,\mathcal{A}(x))$. Note that for any $j\in J$, the stable manifold $\mathcal{W}^{s}(p_{j}(x))$ with respect to the stationary point $p_{j}(x)$ continuously deforms in $x$ (in the sense that $\cup_{x\in\mathcal{X}\setminus\widetilde{\mathcal{X}}}({x}\times W^{s}(p_{j}(x)))$ is an injectively immersed submanifold of $\mathcal{X}\setminus\widetilde{\mathcal{X}}\times\mathbb{R}^{m}$). It is not hard to see that $D(x,b)$ is upper semi-continuous in $x$. Moreover, by the Lipschitz property of $D(x,b)$ with respect to $b$, it follows easily that $D(x,b)$ is upper semi-continuous in $(x,b)$. This implies that $\{(x,b):I_{\mathcal{A}(x)}(y_{0},b)\geq\alpha\}=\{(x,b):D(x,b)=0\}$ is measurable for any $0<\alpha\leq 1$. Note that $\{(x,b):I_{\mathcal{A}(x)}(y_{0},b)\geq\alpha\}$ is the entire domain for any $\alpha\leq 0$ and $\{(x,b):I_{\mathcal{A}(x)}(y_{0},b)\geq\alpha\}$ is empty set for any $\alpha>1$. We find that $I_{\mathcal{A}(x)}(y_{0},b)$ is a measurable function with respect to $(x,b)$.
By design, $\mathcal{A}(x)$ is a union of at most countably many manifolds of dimension strictly less than $m$, and therefore has Lebesgue measure zero in $\mathbb{R}^{m}$. If we uniformly choose $b$ from $\nu,\nu]^{m}$, then it is with probability 0 that $y_{0}+b$ lies in the measure zero set $\mathcal{A}(x)$ for any $x$. Thus, we have the following holds for any $x$:
$\int_{[-\nu,\nu]^{m}}I_{\mathcal{A}(x)}(y_{0},b)db=0.$
It follows
$\int_{\mathcal{X}}\int_{[-\nu,\nu]^{m}}I_{\mathcal{A}(x)}(y_{0},b)dbdx=0.$
By Fubini theorem, we obtain
$\int_{[-\nu,\nu]^{m}}\int_{\mathcal{X}}I_{\mathcal{A}(x)}(y_{0},b)dxdb=\int_{\mathcal{X}}\int_{[-\nu,\nu]^{m}}I_{\mathcal{A}(x)}(y_{0},b)dbdx=0.$ (B.1)
If the claim is wrong, i.e., there exists $U\subset[-\nu,\nu]^{m}$ with positive measure such that
$\int_{\mathcal{X}}I_{\mathcal{A}(x,b)}(y_{0})dx>0$
for any $b\in U$. Then
$\int_{[-\nu,\nu]^{m}}\int_{\mathcal{X}}I_{\mathcal{A}(x)}(y_{0},b)dxdb\geq\int_{U}\int_{\mathcal{X}}I_{\mathcal{A}(x)}(y_{0},b)dxdb>0.$
This contradicts (B.1). Thus, this claim is proven.
Remark B.1.
Theorem B.1 and Lee et al. (2016, Theorem 4.8) both rely on the same technique. Lee et al. (2016, Theorem 4.8) states that if a $C^{2}$ function has the property that the Hessian at every saddle point admits a negative eigenvalue, then gradient descent with a fixed step size less than $1/L$ almost surely converges to a local minimizer, whenever it converges, where $L$ is the Lipschitz smoothness constant of this function.
The core geometric idea in both results is that the stable manifolds associated with non-minimizing stationary points have lower-dimension and thus zero measure. In Lee et al. (2016, Theorem 4.8), they use this idea to argue that if the initialization avoids these lower-dimensional stable manifolds, then the sequence generated by gradient descent method converges to a local minimizer. In Theorem B.1, we instead leverage this to show that discontinuities in the correspondence-driven solution $y^{\text{cd}}(x)$ can only occur when the initialization $y_{0}$ lies in the lower-dimensional stable manifolds. As a result, both results make use of perturbations to the initial point to ensure desirable behavior with probability one.
@@ Appendix C Detailed proof
In all the proofs, measure and measurability are understood in the sense of Lebesgue measure.
@@
Let $f:\mathbb{R}^{n}\to\mathbb{R}^{m}$ be $C^{k}$, where $k\geq\max\{n-m+1,1\}$. Let $X\subset\mathbb{R}^{n}$ denote the critical set of $f$, which is the set of points $x\in\mathbb{R}^{n}$ at which the Jacobian matrix of $f$ has rank <m<m. Then the critical value set, that is, the image $f(X)$, has Lebesgue measure 0 in $\mathbb{R}^{m}$.
Detailed proof of Theorem 2.1: Define the indicator function
$I(g,x)=\begin{cases}1\ &g(x,\cdot)\text{ is not Morse}\\ 0\ &g(x,\cdot)\text{ is Morse}.\end{cases}$
See $I(\widetilde{g}_{a},x)$ as a function of both $a$ and $x$, where $\widetilde{g}_{a}(x,y)=g(x,y)+a^{\top}y$. We first prove $I(\widetilde{g}_{a},x)$ is a measurable function with respect to $(x,a)$. Since the set $\{(x,y,a):\nabla_{y}\widetilde{g}_{a}(x,y)=0\text{ and }\det(\nabla_{yy}^{2}\widetilde{g}_{a}(x,y))=0\}$ is the zero set of some continuous functions, it is a Borel set. Note that $\{(x,a):I(\widetilde{g}_{a},x)\geq 1\}=\mathsf{proj}_{(x,a)}\{(x,y,a):\nabla_{y}\widetilde{g}_{a}(x,y)=0\text{ and }\det(\nabla_{yy}^{2}\widetilde{g}_{a}(x,y))=0\}$, where $\mathsf{proj}_{(x,a)}:\mathbb{R}^{n+2m}\to\mathbb{R}^{n+m}$ denotes the projection onto the $(x,a)$ coordinates, then $\{(x,a):I(\widetilde{g}_{a},x)\geq 1\}$ is an analytic set, thus measurable. It follows that $\{(x,a):I(\widetilde{g}_{a},x)\geq\alpha\}=\{(x,a):I(\widetilde{g}_{a},x)\geq 1\}$ is measurable for any $0<\alpha\leq 1$. Note that $\{(x,a):I(\widetilde{g}_{a},x)\geq\alpha\}$ is the entire domain for any $\alpha\leq 0$ and $\{(x,a):I(\widetilde{g}_{a},x)\geq\alpha\}$ is empty set for any $\alpha>1$. We obtain that $I(\widetilde{g}_{a},x)$ is a measurable function with respect to $(x,a)$. By the same argument, if we fix $x$ and see $I(\widetilde{g}_{a},x)$ as a function of $a$ or fix $a$ and see $I(\widetilde{g}_{a},x)$ as a function of $x$, it is also measurable.
Consider the following gradient map
$F:(x,y)\mapsto\nabla_{y}g(x,y).$
Suppose $\Lambda_{x}$ is the critical value set of $F$ at $x$, i.e., if $\lambda\in\Lambda_{x}$ then there exists $y$ such that $F(x,y)=\lambda$ and $\nabla_{y}F(x,y)=\nabla^{2}_{yy}g(x,y)$ degenerate. By Sard’s theorem, the measure of $\Lambda_{x}$ is 0. This implies that, for any fixed $x$, if we uniformly choose $a$ from $\nu,\nu]^{m}$, the probability of $\{-a\in\Lambda_{x}\}$ is 0. Note that $\{-a\in\Lambda_{x}\}$ means $\widetilde{g}_{a}(x,y)$ is not Morse in $y$, i.e., there exists $y$ such that $\nabla_{y}g(x,y)+a=0$ and $\nabla^{2}_{yy}g(x,y)$ degenerate. Therefore, the following holds for any $x$
$\int_{[-\nu,\nu]^{m}}I(\widetilde{g}_{a},x)da=\int_{-\Lambda_{x}\cap[-\nu,\nu]^{m}}1da=0,$
which implies
$\int_{\mathcal{X}}\int_{[-\nu,\nu]^{m}}I(\widetilde{g}_{a},x)dadx=0.$
By Fubini theorem, we can exchange the order of the integral, i.e., the following holds
$\int_{[-\nu,\nu]^{m}}\int_{\mathcal{X}}I(\widetilde{g}_{a},x)dxda=\int_{\mathcal{X}}\int_{[-\nu,\nu]^{m}}I(\widetilde{g}_{a},x)dadx=0.$ (C.1)
If the claim that “We uniformly choose $a$ from $\nu,\nu]^{m}$ and perturb $g(x,y)$ as $\widetilde{g}_{a}(x,y)=g(x,y)+a^{T}y$. Then, with probability one, $\widetilde{\mathcal{X}}$ has measure zero.” is false, i.e., there exists $U\subset[-\nu,\nu]^{m}$ with positive measure such that
$\int_{\mathcal{X}}I(\widetilde{g}_{a},x)dx>0$
for any $a\in U$. Then
$\int_{[-\nu,\nu]^{m}}\int_{\mathcal{X}}I(\widetilde{g}_{a},x)dxda\geq\int_{U}\int_{\mathcal{X}}I(\widetilde{g}_{a},x)dxda>0.$
This contradicts (C.1). So this claim is proven.
@@
$\left|\eta^{\intercal}H(x)\eta\right|$ $=\left|\int_{\mathbb{R}^{n}}\eta^{\intercal}\nabla^{2}_{xx}h_{\xi}(x-z)\eta\phi^{\text{cd}}(z)dz\right|$
$\overset{(i)}{\leq}\overline{f}\int_{\mathbb{R}^{n}}\left|\eta^{\intercal}\nabla^{2}_{xx}h_{\xi}(x-z)\eta\right|dz$
$=\overline{f}\int_{\mathbb{R}^{n}}\left|\eta^{\intercal}\nabla^{2}_{zz}h_{\xi}(z)\eta\right|dz,$
@@
$\nabla_{zz}^{2}h_{\xi}(z)$ $=\nabla_{z}\left(-h_{\xi}(z)\frac{1}{\xi^{2}}h_{\xi}(z)z\right)$
$=h_{\xi}(z)\left(\frac{zz^{\top}}{\xi^{4}}-\frac{I}{\xi^{2}}\right).$
@@
and in $\text{(}ii)$ we use the following numerical result
@@
A subset $S$ of $\mathbb{R}^{n}$ is a semialgebraic set if it is a finite union of sets defined by polynomial equalities of the form $\{(x_{1},..,x_{n})\in\mathbb{R}^{n}:P(x_{1},...,x_{n})=0\}$ and of sets defined by polynomial inequalities of the form $\{(x_{1},..,x_{n})\in\mathbb{R}^{n}:P(x_{1},...,x_{n})>0\}$.
@@
By Definition 2.3, $\widetilde{\mathcal{X}}$ is the set of $x$ such that $g(x,\cdot)$ is not Morse. It is not hard to see $\widetilde{\mathcal{X}}=\text{Proj}_{x}\{(x,y):\nabla_{y}g(x,y)=0,\ \det(\nabla^{2}_{yy}g(x,y))=0\}$, where $\mathsf{proj}_{x}:\mathbb{R}^{n+m}\to\mathbb{R}^{n}$ denotes the projection onto the $x$ coordinates. Note that the partial derivatives of a semi-algebraic function are also semi-algebraic functions (Coste, 2000, Exercise 2.10); the product of two semi-algebraic functions is again a semi-algebraic function; and the zero set of a semi-algebraic function is a semi-algebraic set. It follows that $\{(x,y):\nabla_{y}g(x,y)=0,\ \det(\nabla^{2}_{yy}g(x,y))=0\}$ is a semi-algebraic set. By Tarski-Seidenberg theorem (Bochnak et al., 1992, Theorem 2.2.1), $\widetilde{\mathcal{X}}$ is also a semi-algebraic set. By Coste (2000, Corollary 3.8), every semi-algebraic set has a stratified manifold structure:
$\widetilde{\mathcal{X}}=\bigcup_{i=1}^{N}M_{i},$
and each $M_{i}$ is diffeomorphic to an open hypercube $(0,1)^{d_{i}}$ with dimension $d_{i}$ (hypercube with dimension 0 is a point). Since the measure of $\widetilde{\mathcal{X}}$ is 0, the dimensions of these manifolds are less than or equal n1n-1. According to Falconer (1990, Section 3.2), each submanifold of $\mathbb{R}^{n}$ with topological dimension $m$ has Minkowski dimension $m$, and the Minkowski dimension is finitely stable, i.e.
$\text{dim}_{\mathrm{box}}(\widetilde{\mathcal{X}})=\max_{i=1,...,N}\{\text{dim}_{\mathrm{box}}(M_{i})\}.$
Therefore we conclude that $\text{dim}_{\mathrm{box}}(\widetilde{\mathcal{X}})\leq n-1$.
@@
By definition of Minkowski dimension, we have

### v3

@@ 6 Conclusion
Building on these results, we designed a biased projected SGD-based algorithm SCiNBiO with a cubic-regularized Newton lower-level solver, and provided convergence guarantees together with oracle complexity bounds for the upper-level. Under the additional assumption that all degenerate stationary points are fold bifurcation points, we further obtained the lower-level oracle complexity of SCiNBiO.
Through numerical experiments, we demonstrated the effectiveness of SCiNBiO in both nonconvex-nonconcave minimax problems and bilevel hyperparameter optimization tasks. In the nonconvex-nonconcave minimax setting, SCiNBiO consistently converged from all random initializations, unlike GDA, which exhibited cyclic behaviors and failed to converge in several cases. For bilevel hyperparameter optimization, SCiNBiO outperformed V-PBGD and BOME in terms of both convergence speed and stability, demonstrating its superior performance in solving bilevel optimization problems with nonconvex lower-level objectives.
Our work develops new modeling and algorithmic approaches for bilevel optimization with nonconvex lower-levels under minimal assumptions, offering practical solution methods together with theoretical insights and provable performance guarantees. Future work will aim to extend these results to settings with stochastic objectives or different lower-level solvers.
@@
Appendix A Full Results of Numerical Example
In this section, we present the plots of the correspondence-driven hyperfunction $\phi^{\mathrm{cd}}(x)$ for the remaining random seeds of the numerical example in Section 5.1, namely, seeds 01414 excluding $5$, $10$, $12$, and $13$.
@@ Appendix B Measure-Zero Discontinuities of Correspondence-Driven Solutions under Gradient Descent
Suppose $g(x,y)$ is a $C^{2}$ function such that it is coercive and semi-algebraic in $y$ for any $x$, $\mathcal{M}$ is gradient descent method, and $0<\eta<1/\overline{L}_{g}$. We uniformly choose $b$ from $\nu,\nu]^{m}$, where $\nu>0$ can be any constant, and perturb the initial point $y_{0}$ by setting it to $y_{0}+b$. Then, with probability one, the set of points in $\mathcal{X}\setminus\widetilde{\mathcal{X}}$ at which $y^{\text{cd}}(x)$ is discontinuous has measure zero, where $y^{\mathrm{cd}}(x)$ is defined under the perturbed function $\widetilde{g}(x,y)$.
@@
    Consider the map $F_{x}:y\mapsto y-\eta\nabla_{y}g(x,y)$, then $F_{x}(y)$ is a diffeomorphism on $\mathbb{R}^{m}$;
@@
Next, we prove that $F_{x}$ is surjective. For any $y_{1}$, consider the following problem
@@
Detailed proof of Theorem B.1: We may assume without loss of generality that $\mathcal{X}\setminus\widetilde{\mathcal{X}}$ is connected; otherwise, since it has at most countably many connected components, the following argument can be repeated on each component and the conclusions then follow by a countable union. Since $g(x,y)$ is coercive and semi-algebraic in $y$, the gradient descent method with step size $\eta$ generates a bounded sequence that converges to a stationary point for any initialization (Attouch et al., 2013). For any fixed $x\in\mathcal{X}\setminus\widetilde{\mathcal{X}}$, $g(x,y)$ is a Morse function; thus, its stationary points are non-degenerate and isolated. Due to coercivity, the set of stationary points is bounded and therefore finite. We denote these points by $\{p_{j}(x)\}_{j\in J}$.
The basin of attraction for each stationary point $p_{j}(x)$ corresponds to its stable manifold $\mathcal{W}^{s}(p_{j}(x))$. According to the Global Stable Manifold Theorem, the stable manifold of a local minimizer is an open $m$-dimensional submanifold, whereas the stable manifolds of non-minimizing stationary points (saddle points and local maximizers) have dimension strictly less than $m$.
Discontinuities in $y^{\text{cd}}(x)$ arise only if the initialization $y_{0}$ lies in the stable manifold of a non-minimizing stationary point. Let $J^{\prime}\subset J$ denote the indices corresponding to such non-minimizers. We define the critical set of initializations as $\mathcal{A}(x):=\bigcup_{j\in J^{\prime}}\mathcal{W}^{s}(p_{j}(x))$. We introduce the indicator function for the perturbed initialization falling into this set:
$I_{\mathcal{A}}(x,b)=\mathbb{I}\{y_{0}+b\in\mathcal{A}(x)\}.$
To apply Fubini’s theorem, we must establish that $I_{\mathcal{A}}(x,b)$ is a measurable function. We define the graph of the critical set in the product space as $\mathcal{G}=\{(x,y):y\in\mathcal{A}(x)\}$. A point $(x,y)$ belongs to $\mathcal{G}$ if and only if the gradient descent sequence initiated at $y$ under the parameter $x$ converges to some non-minimizing stationary point $p_{j}(x)$ for $j\in J^{\prime}$. Let $\Phi_{k}(x,y)$ denote the $k$-th iterate of gradient descent. Since $g$ is $C^{2}$ and $\nabla_{y}g$ is $C^{1}$, $\Phi_{k}(x,y)$ is continuous jointly in $(x,y)$. Furthermore, by the implicit function theorem, the stationary points $p_{j}(x)$ depend continuously on $x$. The convergence condition can be expressed using the Borel hierarchy:
$\mathcal{G}=\bigcup_{j\in J^{\prime}}\bigcap_{m=1}^{\infty}\bigcup_{N=1}^{\infty}\bigcap_{k=N}^{\infty}\left\{(x,y)\in(\mathcal{X}\setminus\widetilde{\mathcal{X}})\times\mathbb{R}^{m}:\|\Phi_{k}(x,y)-p_{j}(x)\|<\frac{1}{m}\right\}.$
The set inside the braces is the inverse image of an open interval under a continuous function, making it an open set in the product space. Since $\mathcal{G}$ is formed by countable unions and intersections of open sets, it is a Borel set. Consequently, the set $\mathcal{S}=\{(x,b):(x,y_{0}+b)\in\mathcal{G}\}$ is also a Borel set, which implies that the indicator function $I_{\mathcal{A}}(x,b)$ is Borel measurable (and thus Lebesgue measurable).
For any fixed $x$, $\mathcal{A}(x)$ is a finite union of submanifolds with dimension less than $m$, so it has Lebesgue measure zero in $\mathbb{R}^{m}$. Therefore, integrating with respect to $b$:
$\int_{[-\nu,\nu]^{m}}I_{\mathcal{A}}(x,b)\,db=0,\quad\forall x\in\mathcal{X}\setminus\widetilde{\mathcal{X}}$.
By Fubini’s theorem, we can swap the order of integration:
$\int_{[-\nu,\nu]^{m}}\left(\int_{\mathcal{X}\setminus\widetilde{\mathcal{X}}}I_{\mathcal{A}}(x,b)\,dx\right)db=\int_{\mathcal{X}\setminus\widetilde{\mathcal{X}}}\left(\int_{[-\nu,\nu]^{m}}I_{\mathcal{A}}(x,b)\,db\right)dx=0.$ (B.1)
This implies that for almost every perturbation $b\in[-\nu,\nu]^{m}$, the integral $\int_{\mathcal{X}\setminus\widetilde{\mathcal{X}}}I_{\mathcal{A}}(x,b)\,dx$ is zero. Since the integrand is non-negative, this means the set of $x$ where $y^{\text{cd}}(x)$ is discontinuous has Lebesgue measure zero.
@@ Appendix C Detailed proof
Let $f:\mathbb{R}^{n}\to\mathbb{R}^{m}$ be $C^{k}$, where $k\geq\max\{n-m+1,1\}$. Let $X\subset\mathbb{R}^{n}$ denote the critical set of $f$, which is the set of points $x\in\mathbb{R}^{n}$ at which the Jacobian matrix of $f$ has rank strictly less than $m$. Then the critical value set, that is, the image $f(X)$, has Lebesgue measure 0 in $\mathbb{R}^{m}$.
Detailed proof of Theorem 2.1: Define the indicator function $I(a,x)=\mathbb{I}\{\widetilde{g}_{a}(x,\cdot)\text{ is not Morse}\}$, where $\widetilde{g}_{a}(x,y)=g(x,y)+a^{\top}y$. First, we establish the measurability of $I(a,x)$. The set of points $(x,y,a)$ satisfying $\nabla_{y}\widetilde{g}_{a}=0$ and $\det(\nabla_{yy}^{2}\widetilde{g}_{a})=0$ is the zero set of continuous functions, and thus a Borel set in $\mathbb{R}^{n+2m}$. Its projection onto the $(x,a)$-space is an analytic set, which is Lebesgue measurable. Consequently, $I(a,x)$ is a measurable function. Next, fix $x\in\mathcal{X}$. The function $\widetilde{g}_{a}(x,\cdot)$ fails to be Morse if and only if a-a is a critical value of the map $y\mapsto\nabla_{y}g(x,y)$. By Sard’s theorem, the set of such critical values has Lebesgue measure zero. Therefore, for any $x$, we have
$\int_{[-\nu,\nu]^{m}}I(a,x)\,da=0.$
Integrating over $\mathcal{X}$, we obtain $\int_{\mathcal{X}}\int_{[-\nu,\nu]^{m}}I(a,x)\,da\,dx=0$. By Fubini’s theorem, we can exchange the order of integration:
$\int_{[-\nu,\nu]^{m}}\left(\int_{\mathcal{X}}I(a,x)\,dx\right)da=0.$
Since the inner integral $\int_{\mathcal{X}}I(a,x)\,dx$ is non-negative, it must be zero for almost every $a\in[-\nu,\nu]^{m}$. This proves that with probability one over the random draw of $a$, the set $\widetilde{\mathcal{X}}$ (where $\widetilde{g}_{a}$ is not Morse) has measure zero.
@@
$\left|\eta^{\intercal}H(x)\eta\right|$ $=\left|\int_{\mathbb{R}^{n}}\eta^{\intercal}\nabla^{2}_{xx}h_{\xi}(x-z)\eta\phi^{\text{cd}}(z)dz\right|\overset{(i)}{\leq}\overline{f}\int_{\mathbb{R}^{n}}\left|\eta^{\intercal}\nabla^{2}_{xx}h_{\xi}(x-z)\eta\right|dz=\overline{f}\int_{\mathbb{R}^{n}}\left|\eta^{\intercal}\nabla^{2}_{zz}h_{\xi}(z)\eta\right|dz,$
@@
$\nabla_{zz}^{2}h_{\xi}(z)$ $=\nabla_{z}\left(-h_{\xi}(z)\frac{1}{\xi^{2}}h_{\xi}(z)z\right)=h_{\xi}(z)\left(\frac{zz^{\top}}{\xi^{4}}-\frac{I}{\xi^{2}}\right).$
@@
and in (ii) we use the following numerical result
@@
A subset $S$ of $\mathbb{R}^{n}$ is a semialgebraic set if it is a finite union of sets defined by polynomial equalities of the form $\{x\in\mathbb{R}^{n}:p(x)=0\}$ and of sets defined by polynomial inequalities of the form $\{x\in\mathbb{R}^{n}:p(x)<0\}$.
@@
Lemma C.1 (Yomdin-Gromov Algebraic Lemma (Yomdin, 1987; Gromov, 1987)).
Let $X\subset[0,1]^{n}$ be a semialgebraic set of dimension $\mu$ defined by conditions
$p_{j}(x)=0\quad\text{or}\quad p_{j}(x)<0,$
where $p_{j}$ are polynomials and $\sum\deg p_{j}=\beta$. Let $r\in\mathbb{N}$. There exists a constant $C=C(n,\mu,r,\beta)$ and semialgebraic functions
$\phi_{1},\ldots,\phi_{C}:(0,1)^{\mu}\to X$
such that their images cover $X$ and $\|\phi_{j}\|_{r}\leq 1$ for $j=1,\ldots,C$, where $\|\phi_{j}\|_{r}:=\max_{\alpha\leq r}\|D^{\alpha}\phi_{j}\|/\alpha!$.
Detailed proof of Theorem 4.1: Let $\mathcal{X}^{*}$ denote the set of $x\in\mathcal{D}$ such that the lower-level function $g(x,\cdot)$ is not Morse, where $\mathcal{D}$ is the projection of the natural domain of $g$ onto the $x$-coordinates. By Definition 2.3, $\widetilde{\mathcal{X}}=\mathcal{X}^{*}\cap\mathcal{X}$, where $\mathcal{X}$ is the compact constraint set of upper-level problem from Assumption 3.2(2). It is not hard to see that $\mathcal{X}^{*}=\text{Proj}_{x}\{(x,y):\nabla_{y}g(x,y)=0,\ \det(\nabla^{2}_{yy}g(x,y))=0\}$, where $\text{Proj}_{x}:\mathbb{R}^{n+m}\to\mathbb{R}^{n}$ denotes the projection onto the $x$ coordinates. Note that the partial derivatives of a semi-algebraic function are also semi-algebraic functions (Coste, 2000, Exercise 2.10); the product of two semi-algebraic functions is again a semi-algebraic function; and the zero set of a semi-algebraic function is a semi-algebraic set. It follows that the set $\{(x,y):\nabla_{y}g(x,y)=0,\ \det(\nabla^{2}_{yy}g(x,y))=0\}$ is a semi-algebraic set in $\mathbb{R}^{n+m}$. By the Tarski-Seidenberg theorem (Bochnak et al., 1992, Theorem 2.2.1), the projection of a semi-algebraic set is semi-algebraic. Thus, the projection $X^{*}$ is a semi-algebraic set in $\mathbb{R}^{n}$.
To bound the Minkowski dimension of $\widetilde{\mathcal{X}}$, we construct a bounded semi-algebraic superset. Since the constraint set $\mathcal{X}$ is compact, there exists a sufficiently large radius $R>0$ such that $\mathcal{X}$ is contained in the closed Euclidean ball $B_{R}=\{x\in\mathbb{R}^{n}:\|x\|\leq R\}$. We define the $Z=\mathcal{X}^{*}\cap B_{R}$. Since $B_{R}$ is semi-algebraic and the intersection of semi-algebraic sets remains semi-algebraic, $Z$ is a bounded semi-algebraic set. Furthermore, by construction, we have the inclusion $\widetilde{\mathcal{X}}\subseteq Z$.
Let $\mu$ be the dimension of $Z$. By Assumption 2.1, $\mathcal{X}^{*}$ has Lebesgue measure zero. Since $Z\subset\mathcal{X}^{*}$, it follows that the Lebesgue measure of $Z$ is also zero. For semi-algebraic sets, zero measure implies that the dimension is strictly less than the ambient dimension, i.e., $\mu\leq n-1$.
We apply the Yomdin-Gromov Algebraic (Lemma C.1) to $Z$. This lemma guarantees the existence of finitely many semi-algebraic maps $\phi_{j}:[0,1]^{\mu}\to\mathbb{R}^{n}$ such that $Z\subseteq\bigcup_{j}\text{Im}(\phi_{j})$ and the first-order derivatives of each $\phi_{j}$ are uniformly bounded. The boundedness of derivatives ensures that each $\phi_{j}$ is Lipschitz continuous. It is a standard property that Lipschitz mappings do not increase the upper Minkowski dimension. Since the upper Minkowski dimension of the unit hypercube $0,1]^{\mu}$ is exactly $\mu$, we have $\overline{\dim}_{\mathrm{box}}(\text{Im}(\phi_{j}))\leq\overline{\dim}_{\mathrm{box}}([0,1]^{\mu})=\mu$.
By the finite stability of the upper Minkowski dimension (i.e., $\overline{\dim}_{\mathrm{box}}(\cup_{j}A_{j})=\max_{j}\overline{\dim}_{\mathrm{box}}(A_{j})$) and its monotonicity with respect to inclusion, we obtain:
$\overline{\dim}_{\mathrm{box}}(\widetilde{\mathcal{X}})\leq\overline{\dim}_{\mathrm{box}}(Z)\leq\overline{\dim}_{\mathrm{box}}\left(\bigcup_{j}\text{Im}(\phi_{j})\right)=\max_{j}\overline{\dim}_{\mathrm{box}}(\text{Im}(\phi_{j}))\leq\mu.$
Given that $\mu\leq n-1$, we conclude that $\overline{\dim}_{\mathrm{box}}(\widetilde{\mathcal{X}})\leq n-1$.
@@
By definition of upper Minkowski dimension, we have